OpenStudy (anonymous):

A silver wire 2m long is to have a resistance of 0.5ohm.What should its diameter be?

OpenStudy (shayaan_mustafa):

simple \[\rho(resistivity)=1.59*10^{-8}\Omega m\] this value is for silver just. use the formula, \[R=\rho L/A\] \[A=\rho L/R\] \[\pi(d/2)^{2}=\rho L/A\] \[(d/2)^{2}=\rho L/\pi A\] \[(d/2)=\sqrt{\rho L/\pi A}\] \[d=2\sqrt{\rho L/\pi A}\] now you can calculate diameter by using above formula. Regards, Read my Profile. :D